# Determinant Estimation under Memory Constraints and Neural Scaling Laws

### v1

@@ 4.2 Larger Data Sets
In our final experiment, we evaluate NTK matrices at an unprecedented scale, where exact determinant computation has never been previously reported. Due to memory constraints, these matrices cannot be stored explicitly, making MEMDET essential for obtaining ground truth values.
We consider two large-scale NTK matrices: $\mathbf{K}_{$50000$}$, a dense matrix of size $500000$\times$500000$$, for ResNet50 trained on CIFAR-10, and an identical-sized NTK for ResNet9 trained on CIFAR-10. At this scale, double-precision arithmetic is infeasible due to memory and computational constraints, making 32-bit computation the only viable direct baseline. However, even in 32-bit, these matrices require nearly a terabyte of storage.
As shown in Figure 5, FLODANCE with $q=6$, $n_{0}=100$, and $n_{s}=5000$ achieves an absolute error of just $0.05\%$ for $L_{$50000$}$ on ResNet9, reducing computation time by a factor of $(n/n_{s})^{3}=1000$. Similarly, for ResNet50, FLODANCE with $q=4$, $n_{0}=100$, and $n_{s}=5000$ achieves an absolute error of $0.02\%$ with the same speedup. In contrast, SLQ exhibited a relative error of 55% (see Table 1, also LABEL:list:imate for implementation). Given their poor performance on smaller datasets, pseudo-NTK and block-diagonal approximations are omitted.
This experiment represents the first exact computation of an NTK determinant at this scale, establishing a new benchmark for large-scale log-determinant estimation.
@@ 5 Conclusion
The calculation of large matrix log-determinants is a commonly encountered but often avoided problem when considering statistical and machine learning problems at scale. A number of techniques have previously been proposed to circumvent explicit computation, typically relying upon stochastic approximations. However, in many problems of interest the sheer size of the matrices, combined with their highly ill-conditioned nature, make not only approximation a difficult task, but forming the matrix itself to provide a baseline becomes computationally intractable. We have addressed this problem on two fronts. On the one hand, we defined MEMDET, a hierarchical memory-constrained algorithm for log-determinant computation, with different versions for general, symmetric, and symmetric positive-definite matrices. On the other hand, we derived neural scaling laws for large kernel matrices, and introduced FLODANCE, a procedure for accurate extrapolation of log-determinants from small subsets of the data. The high level of speed and accuracy of our methods opens the door for routine computation of the interpolating information criteria and related diagnostic tools to enable principled model selection within deep learning frameworks (Hodgkinson et al., 2023b, ).
@@
The ability to accurately compute and estimate matrices of this size further provides fascinating insights into the behavior of the NTKs that we considered in our experiments, which considered square matrices of the size up to $500000$. Further, the memory constrained algorithms we considered can be applied to other classes of matrices (Nguyen & Vu, , 2014; Cai et al., , 2015), where they can be expected to unlock similar insights into their scaling behavior. In terms of further computational tools, we hope for the development of blockwise decompositions of large scale Jacobian matrices, in order to circumvent the need to explicitly calculate $\mathbf{J}\mathbf{J}^{\intercal}$.
@@
Acknowledgments. LH is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144). FR was partially supported by the Australian Research Council through an Industrial Transformation Training Centre for Information Resilience (IC200100022). MWM would like to acknowledge the DOE, IARPA, NSF, and ONR for providing partial support of this work.
Appendix A Related Works
@@
Beyond theoretical foundations, numerous works have focused on efficient implementations of block factorizations, particularly for parallel architectures. Stark & Beris, (1992) optimized block LU decomposition for hierarchical distributed memory, aiming to improve data locality while maintaining parallel efficiency. Dongarra et al., (1979) pioneered high-performance implementations of block factorizations, laying the groundwork for modern HPC systems. More recent studies, such as Galoppo et al., (2005) and Barrachina et al., (2008), have extended block LU methods to GPU-based environments, leveraging parallelism but still assuming that intermediate submatrices fit in memory. While these approaches optimize performance in parallel settings, they do not address the challenge of computing factorizations when the full matrix size far exceeds available RAM. Traditional block methods typically assume that at least some intermediate submatrices can reside in memory, whereas our method (MEMDET) explicitly operates under constrained memory settings, utilizing an out-of-core hierarchical block processing approach.
@@
Appendix B Memory and Computation Challenges of NTK Matrices
The empirical NTK serves as a motivating example throughout this work, as it encapsulates many of the computational challenges associated with large-scale matrix operations. Several software packages have been developed to compute NTK Gram matrices for different neural architectures using automatic differentiation frameworks (Novak et al., , 2022; Engel et al., , 2022). However, the full formation of these matrices remains computationally prohibitive, even on commonly used benchmark datasets.
Table B.1 presents the storage requirements for NTK matrices corresponding to various datasets, highlighting their enormous size. For instance, even CIFAR-10 requires terabytes of storage, while ImageNet-1k exceeds exabytes, making full NTK computation infeasible for most practical applications. Despite its theoretical importance, the NTK Gram matrix is rarely used as a practical tool, with approximations often employed to mitigate computational and memory constraints. Minibatching is one common strategy, and batch-wise NTK approximations have been explored for model selection (Immer et al., , 2023), yet extending these estimates to full datasets remains an open challenge. Alternative approximation techniques (Mohamadi et al., , 2023) have been proposed, but their convergence is only guaranteed in spectral norm, limiting their ability to capture the full spectrum of the NTK. In contrast, the log-determinant—a key quantity in this work—encodes information from the entire eigenvalue distribution, making its computation particularly demanding.
Table B.1: Memory requirements (for different data type precisions) to store empirical NTK matrices of common datasets.
@@ Appendix B Memory and Computation Challenges of NTK Matrices
Beyond storage limitations, the computation time for NTK matrices also presents a major challenge. Table B.2 provides estimated compute times for NTK formation across various models and floating-point precisions on an NVIDIA H100 GPU. Even for relatively small datasets like CIFAR-10, NTK computation is expensive, with higher-precision calculations significantly increasing runtime. For large architectures such as ResNet101, double-precision NTK computation can require over a thousand hours, making exact evaluations impractical without algorithmic improvements like those introduced in this work.
Table B.2: Estimated compute time (in hours using NVIDIA H100 GPU) for NTK matrix computation.
@@
Appendix C Implementation of MEMDET Algorithm
The pseudo-code of MEMDET algorithm is given in Algorithm 2 (for generic matrices), Algorithm 3 (for symmetric matrices), and Algorithm 4 (for symmetric positive-definite matrices), which computes the log-determinant of the matrix $\mathbf{M}$. We note that the log-determinant of all sub-matrices $\mathbf{M}_{[:k,:k]}$, or their row/column permutations for $k=1,\dots,m$, can also be readily computed. For instance, for symmetric and positive-definite matrices using Cholesky decomposition $\mathbf{M}=\mathbf{L}\mathbf{L}^{\intercal}$, with lower-triangular matrix $\mathbf{L}$, the logarithm of its positive determinant is $\mathsf{logdet}(\mathbf{M}_{[:k,:k]})=2\sum_{i=1}^{k}\log(L_{ii})$.
@@ 32                   $\mathcal{R}\leftarrow(j,k+1,k+2,\dots,j-2,j-1)$
Appendix D Optimal Sequence of Processing of Blocks
@@
To find such a sequence of edges, we define $L(G)$, the line graph of $G$ (also known as the edge-to-vertex dual), where each vertex of $L(G)$ represents an edge of $G$. Two vertices in $L(G)$ are connected if and only if their corresponding edges in $G$ share a common vertex. Consequently, any Hamiltonian path in $L(G)$ defines an ordered sequence of edges that fulfills our requirement.
@@ E.1 Computational Complexity
Table E.1 provides a detailed breakdown of the computational complexity of MEMDET for generic matrices (second column, using LU decomposition) and symmetric matrices (third column, using LDL or Cholesky decomposition). The operations are categorized into matrix decomposition, solving triangular systems, and matrix multiplications used to form Schur complements. Each operation’s complexity is given in terms of the number of times it is performed and the FLOP count per operation. In this analysis, a single floating-point fused multiply-add (FMA) operation is counted as one FLOP. The table reports a single complexity column for symmetric matrices, covering both LDL and Cholesky decompositions; while LDL involves additional operations such as row permutations and handling the diagonal matrix $\mathbf{D}$, these steps are excluded from the FLOP count, as their computational cost is negligible compared to the dominant terms.
@@ E.2 Data Transfer and Memory Considerations
For $n_{b}\leq 2$, the entire computation can often fit in memory, avoiding disk I/O, requiring no scratchpad space. However, for $n_{b}>2$, the data transfer overhead increases approximately as $\mathcal{O}(n_{b}^{2})$, meaning that choosing an excessively high $n_{b}$ introduces unnecessary I/O costs. Despite this, the hierarchical design of MEMDET efficiently schedules block transfers, mitigating excessive data movement.
@@
Table E.3: Number of concurrent blocks that must be allocated in memory and the total number of blocks allocated on disk during MEMDET execution. The total required memory and disk space are obtained by multiplying the number of allocated blocks by the block size, $b^{2}\beta$ bytes. While the number of concurrent memory-resident blocks remains fixed, the total number of blocks allocated on disk increases with $n_{b}>2$.
@@ E.3 Empirical Performance Evaluation
Figure E.3: Breakdown of MEMDET runtime (using Algorithm 4) into computation and data transfer times, normalized by $m^{3}$, for $m=2^{13},\dots,2^{16}$. The total process time consists of reading from disk to memory (maroon), writing from memory to disk (light tan), and computation (ochre).
E.4 Conclusing Remarks of Performance Analysis
MEMDET maintains the same computational complexity as conventional factorization methods while distributing computations across blocks. The total FLOP count remains unchanged, but increasing $n_{b}$ shifts the workload between operations (i.e., , from matrix decompositions to matrix multiplications). However, increasing $n_{b}$ also increases data transfer overhead, requiring a balance between reducing memory usage and minimizing disk I/O.
The hierarchical scheduling of block operations optimizes memory usage while limiting unnecessary data movement, ensuring that MEMDET remains efficient under constrained memory conditions. By allowing users to specify a memory limit, MEMDET enables the processing of arbitrarily large matrices on systems with any limited memory size.
@@ E.4 Conclusing Remarks of Performance Analysis
Appendix F Background Material for Neural Scaling Laws
The noiseless result from (Li et al., , 2023) that we require is contingent upon the following assumptions:

### v2

@@ 4.2 Larger Data Sets
In our next experiment, we evaluate NTK matrices at an unprecedented scale, where exact determinant computation has not been previously reported. Due to memory constraints, these matrices cannot be stored explicitly, making MEMDET essential for obtaining ground truth values.
We consider two large-scale NTK matrices: $\mathbf{K}_{$50000$}$, a dense matrix of size $500000$\times$500000$$, for ResNet50 trained on CIFAR-10 with $n=$50000$$ and $d=10$, and an identical-sized NTK for ResNet9. At this scale, computing the full matrix in double precision poses a formidable challenge. To our knowledge, this is the first time an NTK matrix of this size—at full precision and over the entire CIFAR-10 dataset—has been computed, with the matrix itself requiring 2 TB of memory (see Section C.1). This computation was carried out on an NVIDIA Grace Hopper GH200 GPU over 244 hours for ResNet50.
Having established this benchmark, we now evaluate the accuracy of FLODANCE at this scale. As shown in Figure 5, FLODANCE with $q=6$, $n_{0}=100$, and $n_{s}=5000$ achieves an absolute error of just $0.2\%$ for ^$\hat{\ell}_{$50000$}$ on ResNet9, reducing computation time by a factor of $(n/n_{s})^{3}=1000$. Similarly, for ResNet50, FLODANCE with $q=4$, $n_{0}=100$, and $n_{s}=5000$ achieves an absolute error of $0.02\%$ with the same speedup. In contrast, SLQ exhibited a relative error of 55% (see Table 1, also LABEL:list:imate for implementation). Given their poor performance on smaller datasets, pseudo-NTK and block-diagonal approximations are omitted.
This experiment represents the first exact computation of an NTK determinant at this scale, establishing a new benchmark for large-scale log-determinant estimation. Additional experiments are provided in Appendix G, with Figure G.1 demonstrating the accuracy of the global FLODANCE fit, Figure G.2 illustrating sensitivity to subsample choice, and Figure G.3 examining sensitivity to subsample size.
@@ 5 Conclusion
The calculation of large matrix log-determinants is a commonly encountered but often avoided problem when considering statistical and machine learning problems at scale. A number of techniques have previously been proposed to circumvent explicit computation, typically relying upon stochastic approximations. However, in many problems of interest the sheer size of the matrices, combined with their highly ill-conditioned nature, make not only approximation a difficult task, but forming the matrix itself to provide a baseline becomes computationally intractable. We have addressed this problem on two fronts. On the one hand, we defined MEMDET, a memory-constrained algorithm for log-determinant computation, with different versions for general, symmetric, and symmetric positive-definite matrices. On the other hand, we derived neural scaling laws for large kernel matrices, and we introduced FLODANCE, a procedure for accurate extrapolation of log-determinants from small subsets of the data. The high level of speed and accuracy of our methods opens the door for routine computation of interpolating information criteria and related diagnostic tools to enable principled model selection within deep learning frameworks (Hodgkinson et al., 2023b, ).
@@
The ability to accurately compute and estimate matrices of this size further provides fascinating insights into the behavior of the NTKs that we considered in our experiments, which treated square matrices of the size up to $500000$. Further, the memory constrained algorithms we described can be applied to other classes of matrices (Nguyen & Vu, , 2014; Cai et al., , 2015), where they can be expected to unlock similar insights into their scaling behavior. In terms of further computational tools, we hope to develop techniques to extract and extrapolate more refined spectral information about large matrices from small sub-blocks. Methods for blockwise decompositions of large scale Jacobian matrices, would also circumvent the need to explicitly calculate $\mathbf{J}\mathbf{J}^{\intercal}$, enabling higher resolution understanding of their behavior.
@@
Acknowledgments. LH is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144). FR is partially supported by the Australian Research Council through the Industrial Transformation Training Centre for Information Resilience (IC200100022). MWM acknowledges partial support from DARPA, DOE, NSF, and ONR.
Appendix A Nomenclature
Table A.1: Common notations used throughout the manuscript.
Context Symbol Description Example Value Dataset $n$ Number of data points (e.g., images) $50000$ for CIFAR-10 $d$ Number of model outputs (e.g., classes, labels) 10 for CIFAR-10 $m$ Size of NTK matrix is $m\times m$ where $m=nd$ $500000$ for CIFAR-10 MEMDET $n_{b}$ Number of row/column blocks (i.e., $n_{b}^{2}$ blocks in total) e.g., 16 $b$ Block size; typically $b\approx m/n_{b}$, for $b\times b$ submatrices e.g., $31250$ $c$ Available memory capacity (in bytes) $\beta$ Precision (in bytes) of a floating-point number FLODANCE $n_{s}$ Number of sampled data points from $n$ e.g., $2000$ $m_{s}$ Size of sampled NTK matrix $m_{s}=n_{s}d$ e.g., $20000$ $n_{0}$ Start of fitting interval $[n_{0},n_{s}]$ e.g., 100 $q$ Truncation order of Laurent series e.g., 3 $c_{0},\nu_{0},...,\nu_{q}$ Regression coefficients SLQ $l$ Lanczos iterations (Krylov subspace size) e.g., $100$ $s$ Number of Monte Carlo samples e.g., 100 Variables $\mathbf{K}_{n}$ NTK matrix with $n$ data points (matrix of size $m$) $\mathbf{K}_{n_{s}}$ Sampled NTK matrix with $n_{s}$ data points (matrix of size $m_{s}$) $\ell_{n}$ Log-determinant of NTK with $n$ data points (matrix of size $m$) ^$\hat{\ell}_{n}$ Estimated log-determinant $L_{n}$ Normalized log-determinant $L_{n}=n^{-1}\ell_{n}$ $\boldsymbol{\theta}$ Vectorization of neural network parameters $p$ Dimension of $\boldsymbol{\theta}$ Functions $\mathsf{pdet}$ Pseudo-determinant (product of nonzero eigenvalues) $\mathsf{logabsdet}$ Natural logarithm of the absolute value of determinant
We use boldface lowercase letters for vectors, boldface upper case letters for matrices, and normal face letters for scalars, including the components of vectors and matrices. Table A.1 summarizes the main symbols and notations used throughout the paper, organized by context.
Appendix B Related Works in Numerical Linear Algebra
@@
Beyond theoretical foundations, numerous works have focused on efficient implementations of block factorizations, particularly for parallel architectures. Stark & Beris, (1992) optimized block LU decomposition for hierarchical distributed memory, aiming to improve data locality while maintaining parallel efficiency. Dongarra et al., (1979) pioneered high-performance implementations of block factorizations, laying the groundwork for modern HPC systems. More recent studies, such as Galoppo et al., (2005) and Barrachina et al., (2008), have extended block LU methods to GPU-based environments, leveraging parallelism but still assuming that intermediate submatrices fit in memory. While these approaches optimize performance in parallel settings, they do not address the challenge of computing factorizations when the full matrix size far exceeds available RAM. Traditional block methods typically assume that at least some intermediate submatrices can reside in memory, whereas our method (MEMDET) explicitly operates under constrained memory settings, using an out-of-core hierarchical block processing approach.
@@
Appendix C Memory and Computation Challenges of NTK Matrices
C.1 Storage and Compute Requirements
The empirical NTK serves as a motivating example throughout this work, as it encapsulates key computational challenges associated with large-scale matrix operations. Several software packages have been developed to compute NTK Gram matrices for various neural architectures using automatic differentiation frameworks (Novak et al., , 2022; Engel et al., , 2022). However, the full formation of these matrices remains computationally prohibitive, even on common benchmark datasets.
Table C.1 presents the storage requirements for NTK matrices corresponding to various datasets, highlighting their enormous size. For instance, even CIFAR-10 requires terabytes of storage, while ImageNet-1k exceeds exabytes, making full NTK computation infeasible for most practical applications. Despite its theoretical importance, the NTK Gram matrix is rarely used as a practical tool, with approximations often employed to mitigate computational and memory constraints. Minibatching is one common strategy, and batch-wise NTK approximations have been explored for model selection (Immer et al., , 2023). Yet, extending these estimates to full datasets remains an open challenge. Alternative approximation techniques (Mohamadi et al., , 2023) have been proposed, but their convergence is only guaranteed in spectral norm, limiting their ability to capture the full spectrum of the NTK. In contrast, the log-determinant—a key quantity in this work—encodes information from the entire eigenvalue distribution, making its computation particularly demanding.
Table C.1: Memory requirements (for various floating-point precisions) to store empirical NTK matrices of common datasets. The memory is computed as $(nd)^{2}\beta$, where $n$ is the training set size (second column), $d$ is the number of classes (third column), and $\beta$ is the number of bytes per floating-point value.
@@ C.1 Storage and Compute Requirements
Beyond storage limitations, the computation time for NTK matrices also presents a major challenge. Table C.2 provides estimated compute times for NTK formation across various models and floating-point precisions on an NVIDIA H100 GPU. Even for relatively small datasets like CIFAR-10, NTK computation is expensive, with higher-precision calculations significantly increasing runtime. For large architectures such as ResNet101, double-precision NTK computation can require over a thousand hours, making exact evaluations impractical without algorithmic improvements like those introduced in this work.
Table C.2: Estimated compute time (in hours using an NVIDIA H100 GPU) for NTK matrix computation.
@@
C.2 Data Precisions in Our Computational Pipeline
Our computations follow a multi-stage pipeline, with each stage involving distinct data precision formats and practical constraints:
  1. 1.
    Model Training. All neural networks (e.g., ResNet9, ResNet50) were trained using 32-bit precision, which is the default and standard practice in most deep learning frameworks such as PyTorch.
  2. 2.
    NTK Matrix Computation. The NTK matrix is computed from the trained model and stored in various precisions (e.g., 16-bit, 32-bit, and 64-bit, from the same pre-trained model). The “precision” of the NTK matrix, as referred to throughout the paper, reflects the compute and storage format at this stage. Due to the high cost of forming these matrices, it is often tempting or necessary to compute and store them in lower precisions. Our low-precision experiments highlight the pitfalls of mixed-precision in these cases as per Section 2.1, regardless of the downstream use case.
  3. 3.
    Log-Determinant Computation. Regardless of how the NTK matrix was computed and stored (16-bit, 32-bit, or 64-bit), all log-determinant computations were performed in 64-bit precision across all methods (e.g., MEMDET, SLQ, FLODANCE). This represents a “best-case” mixed-precision setup.
Since MEMDET entirely eliminates memory requirement barriers, it became practical to perform high-precision computations (e.g., 64-bit in stage 3) even on large matrices—thus mitigating common concerns about the overhead associated with higher-precision formats.
Appendix D Implementation of MEMDET Algorithm
The pseudo-code of the MEMDET algorithm is given in Algorithm 2 (generic matrices), Algorithm 3 (symmetric), and Algorithm 4 (symmetric positive-definite), each computing the log-determinant of a matrix $\mathbf{M}$. The log-determinants of leading principal submatrices of a matrix $\mathbf{M}$ (possibly after permutation, depending on the decomposition) can also be readily computed. For example, for symmetric positive-definite matrices, the Cholesky decomposition $\mathbf{M}=\mathbf{L}\mathbf{L}^{\intercal}$, with lower-triangular $\mathbf{L}$, gives $\mathsf{logdet}(\mathbf{M}_{[:k,:k]})=2\sum_{i=1}^{k}\log(L_{ii})$, where $\mathbf{M}_{[:k,:k]}$ is the $k\times k$ leading principal submatrix.
@@ 32                   $\mathcal{R}\leftarrow(j,k+1,k+2,\dots,j-2,j-1)$
D.1 Optimal Sequence of Processing of Blocks
@@
To find such a sequence of edges, we define $L(G)$, the line graph of $G$ (also called the edge-to-vertex dual), where each vertex of $L(G)$ represents an edge of $G$. Two vertices in $L(G)$ are adjacent if and only if their corresponding edges in $G$ share a vertex. Thus, any Hamiltonian path in $L(G)$ yields an ordered edge sequence fulfilling our requirement.
@@ E.1 Computational Complexity
Table E.1 provides a detailed breakdown of the computational complexity of MEMDET for generic matrices (second column, using LU decomposition) and symmetric matrices (third column, using LDL or Cholesky decomposition). The operations are categorized into matrix decomposition, solving triangular systems, and matrix multiplications used to form Schur complements. Each operation’s complexity is given in terms of the number of times it is performed and the FLOP count per operation. In this analysis, one FLOP refers to a fused multiply-add (FMA) operation—one multiplication and one addition—as counted by modern GPU benchmarks. The table lists a unified complexity column for symmetric matrices, encompassing both LDL and Cholesky decompositions. While LDL includes additional operations such as row permutations and diagonal scaling via $\mathbf{D}$, these are excluded from the FLOP count due to their negligible cost relative to the leading terms.
@@ E.2 Data Transfer and Memory Considerations
For $n_{b}\leq 2$, the entire computation is performed in memory, avoiding disk I/O and thus requiring no scratchpad space. However, for $n_{b}>2$, the computation utilizes scratchpad space, and data transfer overhead increases approximately as $\mathcal{O}(n_{b}^{2})$, meaning that choosing an excessively high $n_{b}$ introduces unnecessary I/O costs. Despite this, the hierarchical design of MEMDET efficiently schedules block transfers, mitigating excessive data movement.
@@
Table E.3: Number of concurrent blocks that must be allocated in memory (first row) and the total number of blocks allocated on disk (second row) during MEMDET execution. The total required memory and disk space are obtained by multiplying the number of allocated blocks by the block size, $b^{2}\beta$ bytes. While the number of concurrent memory-resident blocks remains fixed, the total number of blocks allocated on disk increases with $n_{b}>2$.
@@ E.3 Empirical Performance Evaluation
Figure E.3: Breakdown of MEMDET runtime (using Algorithm 4) into computation and data transfer times, normalized by $m^{3}$, for $m=2^{13},\dots,2^{16}$. The total process time consists of reading from disk to memory (maroon), writing from memory to disk (light tan), and CPU computation (ochre).
E.4 Concluding Remarks of Performance Analysis
MEMDET maintains the same computational complexity as conventional factorization methods while distributing computations across blocks. The total FLOP count remains unchanged, but increasing $n_{b}$ shifts the workload between operations (i.e., from matrix decompositions to matrix multiplications). However, increasing $n_{b}$ also increases data transfer overhead, requiring a balance between reducing memory usage and minimizing disk I/O.
The scheduling design of block operations optimizes memory usage while limiting unnecessary data movement, ensuring that MEMDET remains efficient under constrained memory conditions. By allowing users to specify a memory limit, MEMDET enables the processing of arbitrarily large matrices on systems with any limited memory size.
@@ E.4 Concluding Remarks of Performance Analysis
E.5 Computational Accuracy of MEMDET
We validate the numerical accuracy of MEMDET by comparing its log-determinant output across various precision formats against two standard in-memory methods: numpy.linalg.slogdet and numpy.linalg.eigh (from which the log-determinant is computed as the sum of the logarithms of the eigenvalues). Table E.4 shows that MEMDET matches both methods with relative errors between $10^{-8}$ and $10^{-16}$, well within the margin of numerical agreement between the baselines themselves.
Table E.4: Relative error of MEMDET (using $n_{b}=8$ blocks) and Numpy’s eigh with respect to NumPy’s slogdet (used as baseline) across different models, dataset sizes, and precision formats.
Quantity Model $\rightarrow$ ResNet9 ResNet9 ResNet18 MobileNet
Dataset CIFAR-10 CIFAR-10 CIFAR-10 MNIST
Subsample Size $n=1000$ $n=2500$ $n=1000$ $n=2500$
Rel. Error MEMDET (16-bit) $2.2\times 10^{-15}$ $5.4\times 10^{-15}$ $5.4\times 10^{-14}$ $8.7\times 10^{-16}$
MEMDET (32-bit) $4.2\times 10^{-12}$ $4.3\times 10^{-10}$ $1.7\times 10^{-11}$ ${8.4\times 10^{-12}}$
MEMDET (64-bit) $1.4\times 10^{-8\phantom{0}}$ $5.4\times 10^{-10}$ $1.8\times 10^{-9\phantom{0}}$ $1.7\times 10^{-6\phantom{0}}$
eigh (16-bit) $1.6\times 10^{-14}$ $1.7\times 10^{-14}$ $2.8\times 10^{-14}$ $1.8\times 10^{-14}$
eigh (32-bit) $5.0\times 10^{-10}$ $1.1\times 10^{-9\phantom{0}}$ $9.2\times 10^{-12}$ $4.5\times 10^{-12}$
eigh (64-bit) $6.1\times 10^{-8\phantom{0}}$ $7.5\times 10^{-9\phantom{0}}$ $1.4\times 10^{-9\phantom{0}}$ $1.0\times 10^{-5\phantom{0}}$
To assess behavior at larger scales, we compute log-determinants of growing NTK submatrices derived from ResNet9, with the matrices formed in both 32-bit and 64-bit floating-point formats. Figure E.4 confirms that MEMDET remains in tight agreement with slogdet and eigh across all scales. Notably, discrepancies between the 32-bit and 64-bit curves are attributable solely to differences in the data precision of the underlying input NTK matrices.
Figure E.4: Log-determinants of growing NTK submatrices from ResNet9, computed in both 32-bit (red) and 64-bit (black) formats. Each submatrix has size $m=nd$ with $d=10$. (a) Comparison between MEMDET, slogdet, and eigh for each input matrix. (b) Relative error of MEMDET and eigh with respect to slogdet. Regardless of the matrix precision, all log-determinant computations are performed in 64-bit precision, and errors remain well below $10^{-7}$.
Finally, we evaluate the effect of MEMDET’s block size parameter $n_{b}$, which determines the number of memory partitions used during computation. Figure E.5 shows that even for NTK matrices of size up to $100000$, increasing the number of blocks has no measurable impact on accuracy: all results remain within $10^{-16}$ to $10^{-12}$ relative error compared to the full in-memory LDL decomposition.
Figure E.5: Effect of the number of blocks $n_{b}$ used in MEMDET. (a) Log-determinants computed for NTK submatrices from ResNet50, with different values of $n_{b}$. (b) Relative error with respect to the conventional LDL decomposition ($n_{b}=1$). All results match to within $10^{-16}$ to $10^{-12}$ accuracy, indicating high numerical stability.
Appendix F FLODANCE and Scaling Laws
In this section, we collect supporting theoretical content related to the FLODANCE algorithm. Section F.1 reviews background material on neural scaling laws. Section F.2 provides the proofs of our main lemma and proposition. Finally, Section F.3 presents the derivation of the FLODANCE parameterization.
F.1 Background Material for Neural Scaling Laws
FLODANCE builds on recent theoretical developments in neural scaling laws that characterize generalization error in terms of kernel eigenvalue decay and source smoothness. Here, we review the key assumptions and results from Li et al., (2023) that underpin our bias analysis. These assumptions concern the spectral properties of the kernel, the embedding behavior of the corresponding RKHS, and the regularity of the target function. We restate them below for completeness.
The following result from Li et al., (2023), stated under a noiseless setting consistent with our framework, relies on the following assumptions: